A 10 kg and a 15kg box are touching each other. A 65 N rightward force is applied to the 10 kg box in order to accelerate both boxes across the surface. The surface is frictionless. What is the acceleration of the boxes and the force acting between the boxes (where the blocks are touching)?
1
Expert's answer
2020-11-13T09:53:54-0500
"F=(m_1+m_2)a\\65=(10+15)a\\a=2.6\frac{m}{s^2}\\F'=F-m_1a=65-2.6(10)=39\ N"